Consider the reaction:

A(g)⇌2B(g)

Find the equilibrium partial pressures of A and B for each value of K. Assume that the initial partial pressure of B in each case is 1.0 bar and that the initial partial pressure of Ais 0.0 bar. Make any appropriate simplifying assumptions.

b)K= 1.1×10−4

Pa,Pb= ______bar

c)K= 2.1×105

Pa,Pb= ______bar
Answer #1

b)

Initial Pressure of B = 1.0 bar

At equilibrium, Pressure of A = x and Pressure of B will be (1-2x)

Since the value is too small, it implies that value of x will be roughly equal to 0.5 (making the numerator pretty small)

Solving the equation

Pressure of A = 0.496 bar

Pressure of B = 1 - 2(0.496) = 0.008 bar

c)

Initial Pressure of B = 1.0 bar

At equilibrium, Pressure of A = x and Pressure of B will be (1-2x)

Since the value is too large, it implies that value of x will be very close to zero (making the denominator pretty small)

Solving the equation

Pressure of A = 4.76 * 10^(-6) bar

Pressure of B = 1 - 2(4.76 * 10^(-6)) = 0.9999904 bar (approximately 1 bar)